OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (-1, 4) and is parallel to the line y = 2x - 3?

OpenStudy (anonymous):

Parallel means they have the same slope so all you need is the point slope formula \[y - y{1} = m(x - x{1})\] So plug in x,y, and slope ( represented by m) and solve for y. \[y - 4 = 2(x - (-1))\] Solve for y \[y = 2x + 6\]